derione:

Dora is purchasing a $162,000 home with a 30-year mortgage at 5.15%. What is her monthly principal and interest payment?

Nurali:

The formula for EMI is : \[\frac{ p*r*(1+r)^{n}}{ (1+r)^{n} - 1 }\] Here p = 162000 r =5.15/12/100=0.004292 n = 30 x 12=360 Now putting the values in the formula we get,